Flamo:

Find the length of the hypotenuse of a right triangle whose legs are 8 and 15 units in length.

Flamo:

A^2 + b^2 = C^2

Flamo:

8^2 + 15^2 = c^2

Angle:

yup!

Flamo:

64 + 225 = c^2

Flamo:

289 To the Square root will be 17.

Flamo:

17 is Right?

Angle:

yup!

Flamo:

Ye!